# Initializing Services in Interactive ML Systems for Diverse Users

### v1

@@ 3 Algorithm & Main Results
In this section, we present our initialization algorithm with guarantees, Algorithm 1, and describe how the steps of the algorithm arise naturally in the interactive systems under consideration. Since collecting preference data uniformly across all the $n$ users can be prohibitively expensive, our goal is to get data from a minimal number of users.
@@
  • User behavior (line 5). Given the list of services $\Theta_{t-1}$ users are assumed to choose the best service that minimizes user risk. Users report their losses with respect to the service they choose from the set of existing services $\{\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[n]}$. In practice, this step requires deploying the services and collecting signals of engagement and utility to determine the loss associated with each user, under the behavioral assumption that users are rational agents that choose the best available service. The provider thus needs to measure each user’s loss only in their single chosen service.
@@
  • New Service (line 7-8). Given a selected user $l$, the algorithm queries the preference $\phi_{l}$ of the user and centers the new service at that preference $\theta_{t}=\phi_{l}$. In practice, this step requires acquiring data about the user in order to learn their preference (this is needed for only $k$ total users throughout the algorithm). For example, data may be acquired by incentivizing the selected users, via offering discount coupons or free premium subscriptions (Hirnschall et al., 2018).
@@
Theorem 1.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Let $\Theta_{\rm OPT}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the total loss and $\mathcal{C}(\Theta_{\rm OPT})$ the resulting partitioning of users (Definition 2.1). If Algorithm 1 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
@@
$\max_{\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})}\frac{4}{\min\limits_{j\in\mathcal{B}}\sum\limits_{i\in\mathcal{B}}c_{ij}}\left(\max_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right).$ (2)
A detailed proof is presented in Appendix B; we summarize the main ideas here. The intuition is that a chosen user’s preference is typically a good representative for other users in its cluster. Thus adding a service parameterized by the chosen user’s preferences generally reduces the losses of users in this cluster. Subsequently we are less likely to pick another user from the same cluster. The $\log k$ factor is due to clusters from which users were never picked.
A similar proof approach was used by Arthur and Vassilvitskii (2007) in the context of the $k$-means problem, by sequentially placing centers on known points sampled with probability proportional to the point’s squared distance to its closest existing center. A key novelty of our analysis is to capture the alignment of diverse loss geometries across users in a large class of functions, specifically understanding how user similarities $c_{ij}$ affect the approximation ratio. [^2] In addition, as stated earlier, we tackle the lack of prior information on $\phi_{i}$ and the function form of $\mathcal{L}_{i}$—this challenge is particular to our setting and does not arise in related standard problems of clustering ($k$-means, $k$-mediods) or resource allocation (facility location problems).
Key characteristics of $K_{\rm OPT}$:
The following are essential characteristics of the term $K_{\rm OPT}$.
@@
    The constant $\min_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}c_{ij}$ captures the user whose loss geometry is least similar to the average loss geometry of the cluster they belong to (recall Assumption 2.i).
@@
    The constant $\max_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}$ captures the user whose preference is least similar to the optimal service parameter of the cluster they belong to (recall Assumption 2.ii).
@@
Fair objective.
While minimizing the total loss is beneficial from the provider’s point of view in keeping users satisfied on average, it is undesirable in human-centric applications if the provided services result in unfavorable or harmful outcomes towards some demographic groups.
Suppose the $n$ users come from $m$ unknown different demographic groups ($m$ is typically small, say racial groups, gender). We denote the groups as $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. The fairness objective is defined as the maximum average loss suffered by any group:
$\Phi(\Theta,\mathcal{A})=\max_{i\in[m]}\frac{\mathcal{L}(\Theta,\mathcal{A}_{i})}{|\mathcal{A}_{i}|}.$ (3)
Ghadiri et al. (2021) defined this objective in the context of fair $k$-means where the points and group identities are known and gave a non-constructive proof that if a $c-$approximate solution for $k$-means exists, it is $m\cdot c$-approximate for fair $k$-means. In our setup, where both user preferences and group identities are unknown, the set of services initialized by Algorithm 1 enjoys guarantees on optimizing the fairness objective as stated below.
Theorem 2.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and unknown associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 1 and 2. Suppose these users belong to $m$ unknown demographic groups $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. Let $\Theta_{\rm fair}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the fairness objective $\Phi$ given in (3). If Algorithm 1 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
$\mathbb{E}_{\Theta_{k}}[\Phi(\Theta_{k},\mathcal{A})]\leq m\cdot K_{\rm OPT}\cdot\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}\cdot(2+\log k)\cdot\Phi(\Theta_{\rm fair},\mathcal{A}),$
where the expectation is taken over the randomization of the algorithm and $K_{\rm OPT}$ is defined in (2).
If the group identities are known a priori to the provider, we can modify Algorithm 1 to get rid of the factor $\frac{\max_{i\in[m]}|\mathcal{A}_{i}|}{\min_{i\in[m]}|\mathcal{A}_{i}|}$ in the approximation ratio, as discussed in Appendix C.
@@ 4 Generalization in Linear Predictors
Theorem 3.
Suppose users belong to $N$ subpopulations satisfying Assumptions 3, 4, and 5. Let $\{n_{i}\}_{i\in[N]}$ denote the number of samples per subpopulation. Let $\Theta_{k}$ be the output of Algorithm 1 using finite sample estimates ^$\widehat{\mathcal{L}}_{i}(\cdot,\phi_{i})$ and $\Theta_{\rm OPT}$ be the optimal solution of the expected loss. Then, for any $\epsilon\in(0,1)$, if $n_{i}=\Omega(\frac{\sigma_{i}^{4}\sqrt{N}\log{(2/\delta)}}{\epsilon^{2}})$ for all $i\in[N]$, the following inequality holds with probability at least $1-\delta$:
$\mathbb{E}_{\Theta_{k}}[\mathcal{L}(\Theta_{k},[N])]\leq\frac{1+\epsilon}{1-\epsilon}K_{\rm OPT}(2+\log k)\mathcal{L}(\Theta_{\rm OPT},[N]),$
where $K_{\rm OPT}$ is as defined in (2) and
$c_{ij}=\frac{1}{2}\min\ \left\{\lambda_{\mathrm{min}}(\widehat{\Sigma}_{i},\widehat{\Sigma_{j}}),\frac{1}{\lambda_{\mathrm{min}}(\widehat{\Sigma}_{i},\widehat{\Sigma_{j}})}\right\},$
with
$\widehat{\Sigma}_{i}=\frac{1}{n_{i}}\sum_{l\in[n_{i}]}x_{i}^{l}(x_{i}^{l})^{\top}\quad\text{and}\quad\widehat{\Sigma}_{j}=\frac{1}{n_{j}}\sum_{l\in[n_{j}]}x_{j}^{l}(x_{j}^{l})^{\top}.$
The proof is presented in Appendix D; we provide a brief overview here. We apply the Chernoff bound to the difference between the empirical loss and expected loss. Note that $(i)$ even if the same set of services are provided, the loss minimizing service for the empirical loss may be different from the expected loss for any subpopulation, and $(ii)$ the optimal set of services for the total empirical loss and the total expected loss are different. Handling these carefully, and utilizing Theorem 1 concludes the proof.
Remark 1.
Note that $\widehat{\Sigma}_{i}$ is the empirical feature covariance of subpopulation $i$. The term $c_{ij}$ captures the alignment between two subpopulation’s loss geometry, and here is equal to half of the minimum generalized eigenvalue of the empirical feature covariances of the the respective subpopulations. This quantity is the largest constant satisfying Assumption 2 (cf. Appendix D).
@@ 5 Experiments
We empirically demonstrate [^4] All our code is available at https://anonymous.4open.science/r/InteractiveSystemInitialization-2A33 the benefits of our algorithm on a commute time prediction task based on 2021 US Census data [^5] https://www.census.gov/programs-surveys/acs/data.html. and a semi-synthetic movie recommendation task on the MovieLens10M dataset. In both the tasks, our algorithm starts without access to any data and our goal is to evaluate how well the sequential acquisition of data, subsequent learning and deployment of learnt models performs on the whole dataset.
Census Data.
@@
We consider the task of predicting daily work commute times, based on 2021 US census data from folktables (Ding et al., 2021). We illustrate a potential use case: the provider is a transport authority offering services in the form of personalized podcasts. If the duration of service is similar to the commute time of a user, that user will be able to consume the media while travelling to work. Hence an accurate prediction of the commute time may be useful in providing services tailored to the users.
The data set has $n=7025$ subpopulations defined by the intersection of gender, race and zip code. We use $k$ linear predictors (as services) with private user features (education, income, transportation mode, age etc). Refer to Appendix E for a detailed description of data pre-processing. As described in Section 4 , the user loss is squared prediction error of their commute time. The features and commute times are unknown a priori to the the provider. At time step $t\leq k$, suppose there exists a set of services $\Theta_{t-1}$. The provider observes the losses (squared prediction error as discussed in Section 4) across different subpopulations {^$\{\widehat{\mathcal{L}}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[N]}$. Based on these observed losses, a subpopulation $l$ is sampled with probability ^$\propto\widehat{\mathcal{L}}_{l}(\Theta_{t-1},\phi_{l})$. Upon selecting a subpopulation the provider surveys the population and acquires relevant data for it. Subsequently the provider uses the acquired data to estimate the sub-population preference via ordinary least squares: $\widehat{\phi}_{l}=\min_{\theta\in\mathbb{R}^{d}}\widehat{\mathcal{L}}_{l}(\theta,\phi_{l})$. The list of services is updated $\Theta_{t}=\Theta_{t-1}\cup\widehat{\phi}_{l}$ and the process is repeated until $k$ services are initialized.
Figure 2: Average loss of different demographics with services initialized by Algorithm 1. Note that Algorithm 1 has much lower losses as compared to the baseline for the each of the demographic groups.
Evaluation.
We compare our algorithm against the baseline where $k$ subpopulations are uniformly selected at random and the services initialized as their learnt regression parameters. The subpopulations belong to the following demographic groups [^6] These group labels are taken from the publicly available 2021 US Census data. (i) Based on sex: male or female, (ii) Based on race: White, Black, American Indian , Alaskan Native, Asian, Pacific Islander, Others. As shown in Figure 2, Algorithm 1 reduces the average loss for every demographic group. We also observe a 15-25% and 7-25% improvement on the fair objective—which is defined in (3)—for sex and racial group demographics, respectively, as the number of services varies (see Figure 4 in Appendix E for detailed plots). We emphasize that this result is observed despite the algorithm not knowing the identity of the demographic groups whilst collecting data.
Figure 3: Total excess error $\sum_{i\in[N]}\mathcal{L}_{i}(V,V_{i})$, where $\mathcal{L}_{i}$ is defined in (4), versus the number of services $d$ (i.e., the number of recommendation models) for Algorithm 1 versus the uniform initialization baseline.
Movie Recommendations.
We conduct a semi synthetic experiment based on the widely used Movielens10M data set (Harper and Konstan, 2015) containing 10000054 ratings across 10681 movies by 71567 viewers. We pre-process the data set (cf. Appendix E) and divide viewers into $N=1000$ subpopulations and retain only the top $m=200$ most watched movies. We compare the performance of Algorithm 1—which initializes a set of recommendation models where user preferences are unknown a priori—to a uniform initialization baseline. For subpopulation $i\in[N]$, let the solution to the following optimization problem denote the unknown user and item embeddings:
$(U_{i},V_{i})=\operatorname*{arg\,min}_{(U,V)\in\mathbb{R}^{n_{i}\times d}\times\mathbb{R}^{d\times m}}\|(UV)_{\Omega_{i}}-r_{i}\|_{2}^{2},$

### v2

@@ 3 Algorithm & Main Results
In this section, we present our initialization algorithm with guarantees, Algorithm 1, and describe how the steps of the algorithm arise naturally in the interactive systems under consideration. Since collecting data uniformly across all the $n$ users can be prohibitively expensive, our goal is to get data from a minimal number of users.
@@
  • User behavior (line 5). Given the list of services $\Theta_{t-1}$ users are assumed to choose the best service that minimizes user loss. Users report their losses with respect to the service they choose from the set of existing services $\{\mathcal{L}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[n]}$. In practice, this step requires deploying the services and collecting signals of engagement and utility to determine the loss associated with each user, under the behavioral assumption that users are rational agents that choose the best available service. The provider thus needs to measure each user’s loss only in their single chosen service.
@@
  • New Service (line 7-8). Given a selected user $l$, the algorithm queries the preference $\phi_{l}$ of the user and centers the new service at that preference $\theta_{t}=\phi_{l}$. In practice, this step requires acquiring data about the user in order to learn their preference parameter (this is needed for only $k$ total users throughout the algorithm). For example, data may be acquired by incentivizing the selected users, via offering discount coupons or free premium subscriptions [21].
@@
Theorem 3.1.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 2.1 and 2.2, with bandit access. Let $\Theta_{\rm OPT}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the total loss and $\mathcal{C}(\Theta_{\rm OPT})$ the resulting partitioning of users (Definition 2.3). If Algorithm 1 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
@@
$\max_{\mathcal{B}\in\mathcal{C}(\Theta_{\rm OPT})}\frac{4}{\min\limits_{j\in\mathcal{B}}\sum\limits_{i\in\mathcal{B}}c_{ij}}\left(\max_{j\in\mathcal{B}}\sum_{i\in\mathcal{B}}\frac{1}{c_{ij}}\right).$
A detailed proof is presented in Appendix D; we summarize the main ideas here. The intuition is that a chosen user’s preference parameter is typically a good representative for other users in its cluster. Thus adding a service parameterized by the chosen user’s preferences generally reduces the losses of users in this cluster. Subsequently we are less likely to pick another user from the same cluster. The $\log k$ factor is due to clusters from which users were never picked.
A similar proof approach was used by Arthur and Vassilvitskii [5] in the context of the $k$-means problem, by sequentially placing centers on known points sampled with probability proportional to the point’s squared distance to its closest existing center. A key novelty of our analysis is to capture the alignment of diverse loss geometries across users in a large class of functions, specifically understanding how user similarities $c_{ij}$ affect the approximation ratio. [^1] In addition, as stated earlier, we tackle the lack of prior information on $\phi_{i}$ and the function form of $\mathcal{L}_{i}$—this challenge is particular to our setting and does not arise in related standard problems of clustering ($k$-means, $k$-mediods) or resource allocation (facility location problems).
Key characteristics of $K_{\rm OPT}$: The following are essential characteristics of the term $K_{\rm OPT}$.
@@
    The constant $\min_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}c_{ij}$ captures the user whose loss geometry is least similar to the average loss geometry of the cluster they belong to (recall Assumption 2.2.i).
@@
    The constant $\max_{j\in\mathcal{B}}\frac{1}{|\mathcal{B}|}\sum_{i}\frac{1}{c_{ij}}$ captures the user whose preference is least similar to the optimal service parameter of the cluster they belong to (recall Assumption 2.2.ii).
@@
Fair objective. While minimizing the total loss is beneficial from the provider’s point of view in keeping users satisfied on average, it is undesirable in human-centric applications if the provided services result in unfavorable or harmful outcomes towards some demographic groups.
Suppose the $n$ users come from $m$ different demographic groups ($m$ is typically small, say racial groups, gender). We denote the groups as $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. The fairness objective is defined as the maximum average loss suffered by any group:
$\Phi(\Theta,\mathcal{A})=\max_{i\in[m]}\mathcal{L}(\Theta,\mathcal{A}_{i})/|\mathcal{A}_{i}|.$ (4)
[15] defined this objective in the context of fair $k$-means where the points and group identities are known and gave a non-constructive proof that if a $c-$approximate solution for $k$-means exists, it is $m\cdot c$-approximate for fair $k$-means. For the fairness objective, we slightly modify our algorithm, by simply reweighting the probability to select an user by the inverse of the size of their demographic group to result in Fair AcQUIre (Algorithm 2 in Appendix E).
Theorem 3.2.
Consider $n$ users with unknown preferences $\{\phi_{1},\ldots,\phi_{n}\}\subset\mathbb{R}^{d}$, and associated loss functions $\mathcal{L}_{i}(\cdot,\cdot)$ satisfying Assumptions 2.1 and 2.2 with bandit access. Suppose these users belong to $m$ demographic groups $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{m}\}\subset[n]$. Let $\Theta_{\rm fair}\subset\mathbb{R}^{d}$ be the set of $k$ services minimizing the fairness objective $\Phi$ given in (4). If Algorithm 2 is used to obtain $k$ services $\Theta_{k}$, then the following bound holds:
@@ 4 Generalization in Linear Predictors
Theorem 4.4.
Suppose users belong to $N$ subpopulations satisfying Assumptions 4.1, 4.2, and 4.3. Let $\{n_{i}\}_{i\in[N]}$ denote the number of samples per subpopulation. Let $\Theta_{k}$ be the output of Algorithm 1 using finite sample estimates ^$\widehat{\mathcal{L}}_{i}(\cdot,\phi_{i})$ and $\Theta_{\rm OPT}$ be the optimal solution of the expected loss. Then, for any $\epsilon\in(0,1)$, if $n_{i}=\Omega(\frac{\sigma_{i}^{4}\sqrt{N}\log{(2/\delta)}}{\epsilon^{2}})$ for all $i\in[N]$, the following inequality holds with probability at least $1-\delta$:
The proof is presented in Appendix F; we provide a brief overview here. We apply the Chernoff bound to the difference between the empirical loss and expected loss. Note that $(i)$ even if the same set of services are provided, the loss minimizing service for the empirical loss may be different from the expected loss for any subpopulation, and $(ii)$ the optimal set of services for the total empirical loss and the total expected loss are different. Handling these carefully, and utilizing Theorem 3.1 concludes the proof.
Remark 4.5.
Note that $\widehat{\Sigma}_{i}$ is the empirical feature covariance of subpopulation $i$. The term $c_{ij}$ captures the alignment between two subpopulation’s loss geometry, and here is equal to half of the minimum generalized eigenvalue of the empirical feature covariances of the the respective subpopulations. This quantity is the largest constant satisfying Assumption 2.2 (cf. Appendix F).
@@ 5 Experiments
(a) Avg. loss (all Census groups)
(b) Avg. loss (worst demographic)
(c) Avg. excess error (ML10M task)
Figure 1: Fig. 1(a) and1(b) show the performance of various user selection strategies on the travel time prediction task on the Census data. Notably, our findings reveal that the greedy and epsilon-greedy baselines exhibit strong performance for $k<10$. However, as the value of $k$ grows, these strategies prove myopic, with random sampling surpassing their effectiveness. AcQUIre and Fair AcQUIre consistently emerge as the two best baselines for both tasks. Fig. 1(c) presents the average excess error for the movie recommendation task. Remarkably, the greedy algorithm demonstrates efficacy when $k$ is small. Epsilon-greedy, employing an explore-vs-exploit approach, successfully overcomes myopic tendencies. Nevertheless, AcQUIre continues to be the best baseline for data collection.
@@
We empirically demonstrate [^3] Code is available at https://anonymous.4open.science/r/MultiServiceInitialization-A422 the benefits of our algorithm on a commute time prediction task based on 2021 US Census data [^4] https://www.census.gov/programs-surveys/acs/data.html. and a semi-synthetic movie recommendation task on the MovieLens10M dataset. In each task the multi-service provider has initially no access to data. Our goal is to evaluate the effectiveness of the iterative data collection and service initialization procedure in comparison to established baselines. Below we first describe both the tasks and then discuss the baselines we consider for our evaluation.
@@
Census Data. We consider the task of predicting daily work commute times, based on 2021 US census data from folktables [13]. We illustrate a potential use case: the provider is a transport authority offering services in the form of personalized podcasts. If the duration of a service is similar to the commute time of a user, that user will be able to consume the media while travelling to work. Hence an accurate prediction of the commute time may be useful in providing services tailored to the users.
The dataset has $N=7025$ subpopulations defined by area zip codes. We use $k$ linear predictors (as services) with user features (education, income, transportation mode, age etc). Refer to Appendix G for details on data pre-processing. The features and commute times are unknown a priori to the provider. At time step $t\leq k$, suppose that the provider offers a set of services parameterized by $\Theta_{t-1}$. The provider observes the losses (squared prediction error of their commute times as discussed in Section 4) across different subpopulations {^$\{\widehat{\mathcal{L}}_{i}(\Theta_{t-1},\phi_{i})\}_{i\in[N]}$. Then the provider selects a subpopulation $l\in[N]$, to get the user feature and commute time data and then estimates $\phi_{l}$ via least squares regression. This selection can be done via our proposed method or one of the baseline strategies discussed later. The provider then centers its next service on $\phi_{l}$ and updates the list of offered services $\Theta_{t}=\Theta_{t-1}\cup\phi_{l}$. Note that in this process the provider only observes features and commute times of users in the $k$ selected subpopulations and $k\ll N$. In Figure 2 we compare runtimes, and observe that even with 1 billion users, AcQUIre takes 300 sec, whereas the greedy and epsilon greedy methods take $>10^{5}$ sec even for 10 million users. With 5000 services, AcQUIre takes $<900$ secs, whereas the runtimes for greedy and epsilon greedy are in the range of $10^{5}$ secs.
Figure 2: Runtimes for AcQUIre and baselines as number of users ($N$) and services ($K$) vary.
Movie Recommendations. We conduct a semi synthetic experiment based on the widely used Movielens10M data set [19] containing 10000054 ratings across 10681 movies by 71567 viewers. We hold out the top $m=200$ movies and pre-process the data set to divide viewers into $N=1000$ user subpopulations based on similarity of their ratings on the remaining movies (cf. Appendix G). Our goal is to evaluate the generalization performance of the baselines to viewers that the provider has never interacted with before. Thus during the service initialization phase, only half of the users in each of the $N$ subpopulations interact with the provider (train set), and we evaluate the performance of the algorithm by the loss incurred by the initialized services on data of the other half of the users that no prior interaction with the provider (test set). For subpopulation $i\in[N]$, the solution to the following optimization problem denotes the user and item embedding: